# SlideSLAM: Sparse, Lightweight, Decentralized Metric-Semantic SLAM for Multi-Robot Navigation

### v5

@@ VI Conclusion
In this work, we introduce a real-time decentralized metric-semantic SLAM framework, along with its integration into autonomous exploration and navigation systems for aerial and ground robots. This integrated system enables teams of heterogeneous robots to autonomously explore and collaboratively construct maps with both geometric and open-vocabulary semantic information across a variety of indoor and outdoor environments. Each robot independently executes the metric-semantic SLAM framework and the autonomous exploration stack onboard in real time, operating in a decentralized manner. The robots opportunistically leverage communication to exchange sparse, lightweight semantic measurements, which are used for inter-robot localization and map merging. Through a comprehensive set of real-world experiments involving three types of aerial and ground robots, we demonstrate the capabilities, effectiveness, and robustness of our system. Quantitative metrics, such as inter-robot localization errors and object mapping accuracy, along with additional results from experiments on the Semantic KITTI dataset, demonstrate the performance of the proposed framework. Our runtime and communication analysis also shows the importance of using sparse semantic map representations, especially for large-scale metric-semantic mapping tasks in challenging and dense environments. For instance, the entire database containing poses and object observations from a mission covering a 1-km trajectory with 1,000 object landmarks amounts to merely 2-3 megabytes. There are several directions for future work. The first is leveraging the heterogeneity in the robot team for active collaborative localization and mapping, building upon our previous works that actively position LiDAR-equipped UGVs as mobile beacons to improve the state estimation [66] or facilitate inter-robot map merging [67] of vision-driven UAVs. The second direction is combining our framework with advanced multi-robot communication frameworks [68] or resource-efficient multi-robot coordination methods [69]. The third direction involves utilizing our sparse, semantically meaningful, open-vocabulary map representation to guide real-time multi-robot decision-making in open-world settings, enabling tasks such as collaborative object search and manipulation in large-scale environments without relying on infrastructure.
VII Acknowledgment

### v6

@@ VII Conclusion
In this work, we introduce a real-time decentralized metric-semantic SLAM framework, along with its integration into autonomous exploration and navigation systems for aerial and ground robots. This integrated system enables teams of heterogeneous robots to autonomously explore and collaboratively construct maps with both geometric and open-vocabulary semantic information across a variety of indoor and outdoor environments. The robots opportunistically leverage communication to exchange sparse, lightweight semantic measurements, which are used for inter-robot localization and map merging. Through a comprehensive set of real-world experiments involving three types of aerial and ground robots, as well as benchmarks on publicly available datasets and comparisons against existing methods, we demonstrate the capabilities, effectiveness, and robustness of our system. Our runtime and communication analysis also show the importance of using sparse semantic map representations, especially for large-scale (over 1 km) metric-semantic mapping tasks in challenging and dense environments, such as forests with over 1000 object landmarks.
VIII Acknowledgment